OpenStudy (anonymous):

What is the simplified form of 3 over 2x plus 5 + 21 over 8 x squared plus 26x plus 15 ? 6 time the quantity 2 x plus 5 end quantity over the quantity 2x minus 5 end quantity times 4 x plus 3 6 over the quantity 4 x plus 3 6 time the quantity x plus 1 end quantity over the quantity 2x minus 5 end quantity times 4 x plus 3 6 time the quantity x plus 1 end quantity over the quantity 2x plus 5 end quantity times 4 x plus 3

OpenStudy (anonymous):

@Nurali

OpenStudy (anonymous):

thanks man

OpenStudy (nurali):

( 3 /( 2x+5 )) + ( 21 / (8x^2 + 26x + 15)) = ( 3 /( 2x+5 )) + ( 21 / (8x^2 + 20x + 6x + 15)) = ( 3 /( 2x+5 )) + ( 21 / (4x(2x + 5) + 3(2x + 5)) = ( 3 /( 2x+5 )) + ( 21 /(2x + 5)(4x + 3) = [ 3 (4x + 3) + 21 ] /(2x + 5)(4x + 3) = [ 12x + 9 + 21 ] /(2x + 5)(4x + 3) = [ 12x + 30 ] /(2x + 5)(4x + 3) = 6(2x + 5) /(2x + 5)(4x + 3) = 6 / (4x + 3)

OpenStudy (nurali):

Anytime.